# Sara Najem

### v1

@@ 5.1 Complexity of the base algorithm
Given adjacency information, computing degrees and forming L=DAL=D-A takes $O(n+m)$ time.
@@
The spectral dimension is estimated from the eigenvalue counting function via repeated linear fits on log-scaled windows. Sorting eigenvalues costs $O(n\log n)$ and scanning candidate windows adds at most another $O(n)$ evaluations, hence $O(n\log n)$ overall (any constant-factor overhead from computing adjusted-$R^{2}$ is dominated by $O(n\log n)$).
@@
For each vertex $i$, curvature is recovered by fitting the diagonal heat kernel values $K(i,i,t)$ to the short-time expansion using a number of sampled time points equal to $\deg(i)$. Let $T$ be the total number of distinct time samples used (across all vertices). Two observations are useful:
  • If $K(t)$ were formed explicitly as a dense matrix for each $t$, computing it would be $O(n^{3})$ per time point, which is unnecessary.
  • Using the eigendecomposition, the diagonal can be computed without forming $K(t)$ explicitly:
    $K(i,i,t)=\sum_{k=1}^{n}V_{ki}^{2}e^{-t\lambda_{k}}.$
    Given $V$ and $\Lambda$, evaluating all $n$ diagonal entries for one $t$ costs $O(n^{2})$ time.
Hence, the total cost of diagonal heat-kernel evaluation is $O(T\,n^{2})$. In our implementation, $T$ is at most linear in the maximum degree (since each vertex uses up to $\deg(i)$ samples), and in particular $T\leq\sum_{i}\deg(i)=2m$ in the extreme case where all vertices use disjoint time grids. Thus, a general upper bound is $O(mn^{2})$ for this stage, although in practice $T$ is much smaller because the time grid is reused across many vertices.
The polynomial fitting step per vertex $i$ uses $O(\deg(i))$ data points. Treating each local regression as $O(\deg(i)^{3})$ yields a worst-case total of $\sum_{i}O(\deg(i)^{3})\leq O(n\Delta^{3})$ where $\Delta$ is the maximum degree. For sparse networks and moderate degrees, this cost is typically dominated by eigendecomposition and/or $O(Tn^{2})$ heat-kernel evaluation.
@@
Putting the pieces together, the worst-case running time of the base algorithm is
@@
In the dense worst case ($m=\Theta(n^{2})$), the dominant term is $O(n^{3})$ from the eigendecomposition together with $O(Tn^{2})$ if $T$ grows with $n$. In the sparse regime with moderate degrees and a shared time grid (small $T$), the eigendecomposition remains the dominant cost, matching the empirical $O(n^{3})$ scaling reported for the implementation.
5.2 Complexity of the generalized algorithm
The generalized algorithm adds two mechanisms: edge subdivisions and vertex identification via triangle attachments. Both preserve equivalence when applied to both graphs, and both trigger re-running the base curvature pipeline.
@@ 5.2 Complexity of the generalized algorithm
Let $G^{(k)}$ denote the graph obtained by subdividing each original edge with $k$ new degree-$2$ vertices. Then the subdivided graph has
$n^{\prime}=n+km\qquad\text{and}\qquad m^{\prime}=(k+1)m.$
Running the base algorithm on $(G^{(k)}_{1},G^{(k)}_{2})$ therefore costs
@@
where $T^{\prime}$ is the number of heat-kernel time samples used after subdivision. The key point is that subdivision preserves polynomial-time behavior in the input size of the transformed instance, but it may substantially increase the constant factors through the growth of $n^{\prime}$.
If subdivision is applied iteratively (increasing $k$ until a spectral smoothness criterion is met), and the algorithm performs $s$ subdivision rounds, then the total work is the sum of the above costs over each intermediate $n^{\prime}$; in the worst case, this is dominated by the final round.
Vertex identification iterations.
When large automorphism groups induce curvature ties, we break symmetry by selecting a vertex from the smallest tied curvature class in each graph and attaching $t$ disjoint triangles in iteration $t$, then re-running the curvature pipeline. At iteration $t$, the number of added vertices is $3t$ (and similarly $3t$ edges). After $r$ iterations, the total added vertices are
@@
so the graph size becomes $n^{\prime\prime}=n^{\prime}+\Theta(r^{2})$.
In the worst case, if we recompute a full eigendecomposition from scratch in every iteration, the total time across $r$ iterations is
@@
which is polynomial in $(n^{\prime}+r^{2})$ but can be large when $r$ grows. Empirically, $r$ remains small in our experiments, so the generalized algorithm is well-approximated by a small constant number of re-runs of the base algorithm on slightly augmented graphs.
@@
Let $n_{\star}$ denote the number of vertices in the largest transformed graph encountered across all subdivision and identification steps. Then the generalized algorithm runs in time
$O\!\left(\#\text{runs}\cdot(n_{\star}^{3}+T_{\star}n_{\star}^{2}+n_{\star}^{2}\log n_{\star})\right),$
where $\text{runs}$ is the number of times the base pipeline is invoked and $T_{\star}$ is the number of time samples used at the largest scale. In particular, for a bounded number of subdivision rounds and identification iterations, the generalized algorithm remains deterministic polynomial time in the size of the transformed instance.
@@
The above bounds assume dense eigendecomposition. In practice, $L$ is sparse for many graph families, and one can replace full diagonalization with iterative methods (e.g., Lanczos) or GPU-accelerated linear algebra to significantly reduce runtime. Such optimizations do not change correctness but can substantially extend the range of feasible $n$ in large-scale evaluations.
@@ 6.4 Regular graphs
We apply our framework on regular graphs of size $N=500$ $\rho$ and $C$ are shown in Figure 9.
@@ 6.9 Paley graphs
We show the results of the symmetry breaking construction for vertex identification in Figure 16. We then compute curvatures associated with the resultant graphs shown in Figure 17. They exhibit a clear breaking of symmetry of the green nodes allowing us to identify an isomorphic mapping between the nodes by the values of sorted their curvatures.
@@ 7 Conclusion and future work
In this work, we introduced a new algorithmic approach to the graph isomorphism problem based on spectral geometry. By extracting curvature information from the Laplacian via heat-kernel-based constructions, we developed a symmetry-breaking mechanism that goes beyond classical spectral invariants. While this method does not constitute a proof that graph isomorphism admits a polynomial-time algorithm, it demonstrates that geometric information derived from spectral data can be surprisingly effective in practice. In extensive experiments on all graph families generated by standard generators, the algorithm correctly resolved every instance tested.
These results suggest that curvature-based spectral features capture structural distinctions that are invisible to eigenvalues alone. More broadly, they reveal a previously unexplored connection between graph isomorphism and discrete spectral geometry, complementing existing combinatorial and group-theoretic approaches. This perspective provides a new lens through which symmetry in graphs can be analyzed and potentially exploited algorithmically.
@@
Second, a deeper theoretical understanding of the method is needed. At present, the success of curvature as a symmetry-breaking tool is supported empirically but not yet fully explained. An important direction is to analyze how curvature interacts with graph automorphism groups, eigenvalue multiplicities, and spectral degeneracies, and to determine under what conditions curvature yields provably unique or canonical vertex distinctions. Establishing formal guarantees, such as bounds on failure probability or connections to known tractable graph classes, would significantly strengthen the theoretical foundations of this approach.
Finally, the ideas developed here are not specific to graph isomorphism. Spectral geometry offers a general framework for extracting geometric structure from discrete objects, and curvature-based features may be useful in a wide range of algorithmic problems where symmetry and structural ambiguity play a central role. Potential applications include graph canonization, clustering, network alignment, and other areas of algorithm design where distinguishing subtle structural differences is critical. Exploring these broader applications may reveal that curvature, much like eigenvalues and eigenvectors before it, becomes a fundamental primitive in the algorithmic toolkit.

### v2

@@ 6.1 Complexity of the base algorithm
Given adjacency information, computing vertex degrees and forming the combinatorial Laplacian L=DAL=D-A takes $O(n+m)$ time.
@@
The spectral dimension is estimated from the eigenvalue counting function via repeated linear fits on logarithmically scaled windows. Sorting eigenvalues costs $O(n\log n)$, and scanning candidate windows adds at most another $O(n)$ operations, yielding $O(n\log n)$ overall. Any constant-factor overhead from computing adjusted-$R^{2}$ values is dominated by this term.
@@
For each vertex $i$, curvature is recovered by fitting the diagonal heat-kernel values $K(i,i,t)$ to the short-time expansion using a number of sampled time points proportional to $\deg(i)$. Let $T$ denote the total number of distinct time samples used across all vertices. Two observations are useful:
  • Explicitly forming $K(t)$ as a dense matrix for each $t$ would cost $O(n^{3})$ per time point, which is unnecessary.
  • Using the eigendecomposition, the diagonal entries can be computed as
    $K(i,i,t)=\sum_{k=1}^{n}V_{ki}^{2}e^{-t\lambda_{k}},$
    so evaluating all $n$ diagonal entries for a single $t$ costs $O(n^{2})$ time.
Thus, diagonal heat-kernel evaluation costs $O(Tn^{2})$ overall. In the worst case, $T\leq\sum_{i}\deg(i)=2m$, yielding an upper bound of $O(mn^{2})$ for this stage, although in practice $T$ is much smaller due to shared time grids.
The polynomial fitting step at vertex $i$ uses $O(\deg(i))$ samples. Treating each local regression as $O(\deg(i)^{3})$ yields a worst-case total of $\sum_{i}O(\deg(i)^{3})\leq O(n\Delta^{3})$, where $\Delta$ is the maximum degree. For sparse graphs with moderate degrees, this cost is typically dominated by eigendecomposition and/or diagonal heat-kernel evaluation.
@@
The worst-case running time of the base algorithm is
@@
In dense graphs ($m=\Theta(n^{2})$), the eigendecomposition term $O(n^{3})$ dominates. In sparse graphs with moderate degrees and shared time grids, the empirical scaling is likewise dominated by $O(n^{3})$.
6.2 Complexity of handling degeneracies
To address spectral degeneracies and large symmetry groups, the algorithm incorporates edge subdivision and vertex identification. Both mechanisms trigger re-execution of the base curvature pipeline.
@@ 6.2 Complexity of handling degeneracies
Let $G^{(k)}$ denote the graph obtained by subdividing each original edge with $k$ new degree-$2$ vertices. The resulting graph has
$n^{\prime}=n+km,\qquad m^{\prime}=(k+1)m.$
Running the base algorithm on $(G^{(k)}_{1},G^{(k)}_{2})$ costs
@@
where $T^{\prime}$ denotes the number of heat-kernel samples used after subdivision. Subdivision preserves polynomial-time behavior in the size of the transformed instance, but may increase constant factors due to the growth of $n^{\prime}$. If subdivision is applied iteratively over $s$ rounds, the total cost is dominated by the final round.
Vertex identification.
When large equivalence classes persist, symmetry is broken by attaching $t$ disjoint triangles to selected vertices in iteration $t$, followed by recomputation. At iteration $t$, $3t$ vertices are added. After $r$ iterations, the total number of added vertices is
@@
so the graph size becomes $n^{\prime\prime}=n^{\prime}+\Theta(r^{2})$. If a full eigendecomposition is recomputed at each iteration, the total cost across $r$ iterations is
@@
which remains polynomial in $(n^{\prime}+r^{2})$. Empirically, $r$ is small, and only a few re-executions are required.
@@
Let $n_{\star}$ denote the number of vertices in the largest transformed graph encountered. The degeneracy-handling stage runs in
$O\!\left(\#\text{runs}\cdot\bigl(n_{\star}^{3}+T_{\star}n_{\star}^{2}+n_{\star}^{2}\log n_{\star}\bigr)\right),$
where $\text{runs}$ is the number of base-pipeline executions and $T_{\star}$ is the maximum number of heat-kernel samples used.
@@
The above bounds assume dense eigendecomposition. In practice, the Laplacian is sparse for many graph families, and iterative eigensolvers (e.g., Lanczos) or GPU-accelerated linear algebra can substantially reduce runtime without affecting correctness.
6.3 Complexity of the full BFS-curvature–signature algorithm
We now consider the complete algorithm, which augments curvature matching with BFS-curvature signatures.
Signature construction.
Assuming all Minakshisundaram–Pleijel coefficients have already been computed, assembling the BFS-curvature signature for a single vertex requires grouping vertices by BFS distance and sorting coefficient within each layer. In the worst case, this costs $O(n\log n)$ per vertex, yielding a total cost of
$O(n^{2}\log n)$
to construct signatures for all vertices.
Signature comparison and refinement.
Sorting and comparing the $n$ BFS-curvature signatures in each graph costs $O(n^{2}\log n)$ in total. If signatures are not unique, degeneracy-handling mechanisms are invoked, triggering recomputation on augmented graphs.
Overall bound.
Let $n_{\star}$ be the size of the largest transformed graph encountered. The total running time of the full algorithm is bounded by
$O\!\left(\#\text{runs}\cdot\bigl(n_{\star}^{3}+T_{\star}n_{\star}^{2}+n_{\star}^{2}\log n_{\star}\bigr)\right).$
The BFS-curvature–signature machinery therefore adds at most quadratic (over logarithmic factors) overhead and does not alter the asymptotic polynomial-time complexity of the method. In all tested instances, spectral decomposition remains the dominant cost.
Summary.
The full BFS-curvature–signature algorithm runs in deterministic polynomial time in the size of the largest transformed instance. Enriching scalar curvature with multi-scale BFS-curvature signatures substantially increases discriminative power while preserving the asymptotic complexity class. The resulting procedure provides a principled local-to-global mechanism for vertex identification and graph isomorphism testing.
@@ 7.4 Regular graphs
We apply our framework on regular graphs of size $N=500$. $\rho$ and $C$ are shown in Figure 9.
@@ 7.9 Paley graphs
We show the results of the symmetry breaking construction for vertex identification in Figure 16. We then compute curvatures associated with the resultant graphs shown in Figure 17. They exhibit a clear breaking of symmetry of the green nodes allowing us to identify an isomorphic mapping between the nodes by the values of their sorted curvatures.
@@ 8 Conclusion and future work
In this work, we introduced a new algorithmic approach to the graph isomorphism problem based on spectral geometry. By extracting curvature information from the graph Laplacian via heat-kernel-based constructions and organizing this information into multi-scale, BFS-based curvature signatures, we developed a principled symmetry-breaking mechanism that goes beyond classical spectral invariants. While this method does not constitute a proof that graph isomorphism admits a polynomial-time algorithm, it demonstrates that geometric information derived from spectral data can be surprisingly effective in practice. In extensive experiments across all graph families generated by standard generators, the algorithm correctly resolved every instance tested.
These results suggest that curvature-based spectral features capture structural distinctions that are invisible to eigenvalues alone. More broadly, they reveal a previously unexplored connection between graph isomorphism and discrete spectral geometry, complementing existing combinatorial, refinement-based, and group-theoretic approaches. Viewed through this lens, vertex identification becomes a local-to-global process: rich local geometric signatures induce a global correspondence between graphs.
@@
Second, a deeper theoretical understanding of the method is needed. At present, the success of curvature-based signatures as symmetry-breaking tools is supported empirically but not yet fully explained. An important direction is to analyze how these signatures interact with graph automorphism groups, eigenvalue multiplicities, and refinement hierarchies, and to determine under what conditions they yield provably unique or canonical vertex distinctions. Establishing formal guarantees, such as sufficient conditions for completeness or connections to known tractable graph classes, would significantly strengthen the theoretical foundations of this approach.
Finally, the ideas developed here are not specific to graph isomorphism. Spectral geometry offers a general framework for extracting geometric structure from discrete objects, and curvature-based features may be useful in a wide range of algorithmic problems where symmetry and structural ambiguity play a central role. Potential applications include graph canonization, clustering, network alignment, and related problems where distinguishing subtle structural differences is critical. Exploring these broader applications may reveal that curvature, much like eigenvalues and eigenvectors before it, becomes a fundamental primitive in the algorithmic toolkit.